OpenStudy (anonymous):

**PLEASE HELP!!!!!!!** Johnathan has 3 times as much money as Christian. Ashley has $4 less than Johnathan and $5 more than Christian. How much do each have?

OpenStudy (anonymous):

You can't determine how much money Ashley is making, because she is a whore, and working the streets right now making money. That's the answer.

OpenStudy (anonymous):

Really I need a serious answer!!

OpenStudy (anonymous):

Ok suzanne for simplicity let us assume Christian as C,Ashley as A and Johnathan as J.Now writing the given statement in forms of mathematical equations J=3C(i.e Johnathan has 3 times the money of Christian) A=J-4 also A=C+5 Solving these three equations will give you the answer i.e 3C-4=C+5 so C=$4.5,thus J=3*4.5=$13.5 and A=$9.5

OpenStudy (anonymous):

thank you soo much!!!!